OpenStudy (anonymous):

PLZZ HELP!! Johnny wants to start making and selling surfboards. His costs are the large sum of money that includes his initial investment and the cost of materials for each surfboard. He has already planned how much he is going to sell his surfboards for. Explain how Johnny can set up a system of equations to predict when he will start to make a profit.

OpenStudy (anonymous):

Hi beautiful :)

OpenStudy (anonymous):

can u not?! I actually need to know how to solve this

OpenStudy (anonymous):

Initial investment = i cost of materials per surfboard = cs price to sell per surfboard =ps

OpenStudy (anonymous):

i + cs = johnny's spending ps = the money coming in. Johnny will start to make a profit when he sells enough surfboards so that his spending equals the money coming in. So, i + cs = ps

OpenStudy (anonymous):

For example, if i=100, c=10, p=20......... 100+10s=20s Solve for s. 100=10s s=10.

OpenStudy (anonymous):

\[I+CS=PS\] is your equation.